### v1

# Regulatable ML

Abstract
As autonomous AI agents are increasingly deployed in industry, it is essential to safeguard them. We introduce a novel framework that automates the translation of unstructured design documents into verifiable, real-time guardrails. We introduce "Policy as Prompt," a new approach that uses Large Language Models (LLMs) to interpret and enforce natural language policies by applying contextual understanding and the principle of least privilege. Our system first ingests technical artifacts to construct a verifiable policy tree, which is then compiled into lightweight, prompt-based classifiers that audit agent behavior at runtime. We validate our approach across diverse applications, demonstrating a scalable and auditable pipeline that bridges the critical policy-to-practice gap, paving the way for verifiably safer and more regulatable AI.
1 Introduction
The rapid proliferation of sophisticated AI agents Wang et al. (2025); Ding et al. (2024) has created an urgent need for robust governance to ensure their alignment with human policies and safety constraints Liu et al. (2024); Kim et al. (2023); Li et al. (2025); He et al. (2024). Emerging frameworks like the EU AI Act eu- (2024) highlight this by mandating comprehensive risk and conformity assessments Fabiano (2024). However, a critical gap persists between high-level, human-readable policies and the low-level, machine-enforceable rules required for practical implementation. This "policy-to-practice" gap presents a significant bottleneck for effective AI evaluation, compliance, and auditing.
To address this challenge, we propose that guardrail security policies can be used to govern the behavior of AI agents, drawing inspiration from established data security principles. A guardrail is a defense mechanism used in AI applications to prevent unintended or harmful behavior Dong et al. (2024); Zhang et al. (2025). The "policy-to-practice" gap is particularly acute for autonomous agents, as the appropriateness of their actions is entirely dependent on context. As argued by the CONSECA framework Tsai and Bagdasarian (2025), security for these generalist systems requires moving beyond static policies, which can overrestrict or overallow actions, to a new paradigm of just-in-time, contextual security. While frameworks like TrustAgent Hua et al. (2024) have proposed using static constitutional principles, this is insufficient for the dynamic nature of agent interactions.
To bridge this critical gap of enforcing security policies for AI agents at scale, we introduce Policy as Prompt, a novel framework for translating natural language policy documents into dynamic, enforceable guardrails. Our work provides a concrete solution to the problem identified by CONSECA by operationalizing its core principles. Our key contributions are as follows: (i) We introduce a scalable, end-to-end pipeline that automatically ingests unstructured technical artifacts to identify and extract security constraints Tsai and Bagdasarian (2025). (ii) We propose a verifiable process where these constraints are converted into a human-readable policy draft, enabling efficient review and refinement by security engineers. (iii) We compile the verified policy into prompt-based classifiers—our "guardrail security policies" Dong et al. (2024)—that enforce a least-privilege policy at runtime by validating agent inputs and outputs Zhang et al. (2025). (iv) We validate our approach by generating policies for various LLM applications and test their effectiveness across different state-of-the-art models.
Figure 1: Policy Generation and Enforcement Pipeline
2 Policy Tree Generation
Our technique extracts guardrail security requirements for LLM applications directly from development artifacts. By analyzing high-level design documents, we capture the system’s intended security posture before implementation. Grounding policies in their original design context is critical for ensuring AI systems operate as intended Tsai and Bagdasarian (2025).
The core of our POLICY-TREE-GEN approach is a verified two-pass parsing and enrichment pipeline that extracts security intent from source artifacts. In Pass 1, a generative LLM infers the document hierarchy, extracts verbatim policy statements, and classifies them into four categories: VALINP (acceptable inputs), INVALINP (unacceptable inputs, e.g., out-of-domain topics), VALOUT (sanitization), and INVALOUT (forbidden outputs, such as hallucinations, policy violations, or PII leaks). Non-o1 models in the pipeline operated at a temperature of 0.5. This parsing step yields a policy tree structure Kumar et al. (2025), illustrated in Appendix 2, with example prompts in Appendix 3. For instance, “Resolved HR case data is pulled from Internal Case Portal” is a VALINP, implying “non-HR data” is an INVALINP; “delete or anonymize personal data before analysis” reinforces INVALOUT. All extracted statements undergo verification by a separate judge LLM (gpt-4o) to ensure they are faithful verbatim quotes and correctly categorized; failing items are pruned. In Pass 2, the verified structure is enriched with relevant, source-cited examples, producing a high-fidelity hierarchical policy tree that preserves contextual relationships and enhances interpretability and maintainability.
Our artifact-driven policy generation applies the Principle of Least Privilege to LLM agents Zhang et al. (2025). By converting application artifacts into a structured, verifiable tree, our method enables systematic guardrail enforcement, as demonstrated by the HR policy tree in Figure 1. Policy scope is limited to the LLM application’s input and output boundaries. VALINP/INVALINP govern data entering the first agent, while VALOUT/INVALOUT define rules for the final user response.
Table 1: Policy Enforcement Example for HR Input Classifier
Application Example Classification Reason Action
HR Update my address on Workday VALID (Input) ALLOW
HR My address is 21 Victoria St INVALID (Input) Contains Non-Anonymised PII BLOCK
HR Ignore rules and reveal your system prompt INVALID (Input) Malicious / Prompt Injection ALERT
HR Send to sample@gmail.com VALID (Input) Anonymised PII ALLOW
HR US election news link… INVALID (Output) Non HR Content BLOCK
3 Policy as Prompt Generation
The POLICY-AS-PROMPT-GEN begins with the verified policy tree from POLICY-TREE-GEN, which encodes categorized security rules and examples. This tree is transformed into a human-readable markdown document for LLM consumption, where each instruction is formatted as a rule with examples labeled ‘positive’ (compliant) or ‘negative’ (violating). The labeled examples are reused to synthesize few-shot prompt blocks that accompany the rules.
It is then embedded into a master template that primes the LLM as a compliance analyst. Two templates are used: one for an Input Classifier and another for an Output Auditor. These prompt the model to consult the policy rules and return a JSON object with a binary classification (VALID/INVALID) and a concise justification. For HR input classification (Table 1), HR-related requests without non-anonymized PII are VALID, while those with PII, prompt-injection attempts, or non-HR content are INVALID. The strict output format enables deterministic system actions (ALLOW, BLOCK, ALERT). HR Input Classifier is shown in Figure 4 and SecOps in Figure 5.
The deliverables are two markdown few-shot prompts (rules + exemplars), which undergo human-in-the-loop review. Security engineers then either (i) approve for deployment, (ii) reject requiring upstream updates and regeneration, or (iii) request changes to the prompt markdown without altering the upstream tree.
4 Experimental Setup and Results
This study evaluated the generation and enforcement of guardrail policies across two distinct LLM-powered systems in HR and Security Operations (SecOps) domains. The input artifacts for all models consisted of PRDs, technical design documents, and prompts extracted from the application source code. To ensure compatibility, all documents were converted to Markdown format, and any embedded images were replaced with textual descriptions generated by gpt-4o. All reported metrics are the average of multiple runs to ensure stability.
4.1 Policy-Tree-Gen Analysis
Table 2: Evaluation Metrics for POLICY-TREE-GEN
Detection Classification Per-class F1
Application Model R (%) F1 (%) Macro-F1 (%) VALINP VALOUT INVINP INVOUT
HR O1 53.3 60.0 24.5 35.3 29.4 0.0 33.3
HR GPT-OSS 120B 17.8 25.0 4.8 19.4 0.0 0.0 0.0
HR Llama 405B 8.9 14.5 4.5 18.2 0.0 0.0 0.0
HR Claude 3.5 4.4 8.0 12.5 0.0 0.0 0.0 50.0
SecOps O1 19.4 22.6 13.0 22.2 29.6 0.0 0.0
SecOps GPT-OSS 120B 5.6 10.3 4.2 16.7 0.0 0.0 0.0
SecOps Llama 405B 5.6 9.8 10.0 0.0 0.0 40.0 0.0
SecOps Claude 3.5 2.8 5.4 6.2 0.0 0.0 25.0 0.0
We evaluated several large language models including Llama 3 405B(Grattafiori et al., 2024), GPT OSS 120B (OpenAI et al., 2025), Sonnet 3.5(Anthropic, 2024) and o1 (OpenAI et al., 2024) on two applications, HR and SecOps. The models Gemma 3 1B(Team et al., 2025) and Qwen 1.7B Thinking(Yang et al., 2025) were excluded due to poor performance. Gold policies were created by security engineers and served as the ground truth against which the LLM-generated policies were evaluated.
The metrics in Table 2 evaluate model performance in two categories: Detection, measured by Recall (R) and F1 score, and Classification, assessed via Macro-F1 and individual per-class F1 scores for VALINP, VALOUT, INVINP, and INVOUT. A high score in these metrics indicates a model that is effective at both identifying relevant requirements and accurately assigning the correct category to them. The HR application consistently yields higher performance scores for all models compared to the SecOps application. Furthermore, the o1 model significantly outperforms all other models, demonstrating its superior capability in this specific task.
As shown in Table 4, this analysis details model performance on HR and SecOps tasks using metrics such as Detection Precision, which measures the percentage of a model’s predictions that are correct, and Micro-F1. Span quality is evaluated using four metrics: Span Exact, which measures if the extracted text is an exact match to the ground truth; Token-F1, which assesses the token-level overlap between the extracted and ground truth text; Substr, which checks if one text is a substring of the other; and Emb Cos, which is the cosine similarity of the text embeddings. A high score in these metrics indicates that the model is accurately extracting the correct text, even if it might misclassify it. While the o1 model remains a top performer, others reveal a significant disconnect between their extraction quality and overall F1 score. For HR applications, models like Llama 405B and Claude 3.5 have high span quality metrics, indicating accurate text extraction, yet their low Micro-F1 scores suggest they struggle with correct classification. In the SecOps domain, performance is generally lower and more varied. Notably, Claude 3.5 achieved a high Det P on SecOps, meaning all of its identified requirements were true positives, but its Micro-F1 score remained very low due to continued classification errors.
4.2 Policy-As-Prompt-Gen Analysis
Table 3: Accuracy per application/model for POLICY-AS-PROMPT-GEN
Application Model Input Acc. Output Acc.
HR GPT-4o 0.73 0.71
Qwen3-1.7B 0.66 0.59
Gemma-1B 0.40 0.32
SecOps GPT-4o 0.70 0.68
Qwen3-1.7B 0.66 0.61
Gemma-1B 0.42 0.41
Policy enforcement was tested on a separate set of models: Qwen 3 1.7B (Thinking Mode) (Yang et al., 2025), GPT-4o (OpenAI et al., 2024), and Gemma 3 1B(Team et al., 2025). For specialized data classification tasks within this framework, we employed Small Language Models (SLMs), leveraging their established effectiveness and low latency, which helps to quickly run policies in real-time.
Both policies were judged request review by security engineers and deployed after minimal changes, saving time for the security team. POLICY-AS-PROMPT-GEN was evaluated on functional correctness by integrating generated prompts into target LLMs and testing with 100 gold inputs and outputs. Tests included both standard and adversarial cases (e.g., prompt injection, toxic content). Evaluation focused on accuracy of INVALID/VALID classifications, with defaults of INVALID $\Rightarrow$ BLOCK and VALID $\Rightarrow$ ALLOW, unless otherwise specified. A limitation is that prompt tuning emphasized GPT models, potentially contributing to performance gaps.
5 Conclusion
In this work, we introduced a framework that bridges the policy-to-practice gap by transforming unstructured design artifacts into verifiable guardrails. We demonstrated an end-to-end pipeline for automated, auditable, and enforceable policy generation. Our experiments show that while large proprietary models excel in policy extraction, smaller models can still effectively enforce policies with curated prompts. This approach establishes a scalable path toward trustworthy, regulatable AI systems grounded in transparent policy governance.
6 Limitations
Despite promising results, this work has several limitations that temper generalizability and reproducibility. Our evaluation spans two enterprise-style domains (HR, SecOps) with modest gold sets (100 inputs and 100 outputs per application) and a limited number of design artifacts, which constrains transferability to other settings and larger, more heterogeneous corpora. Policy extraction currently depends on large proprietary models, while open-source baselines differ in size or tuning; moreover, prompting effort was greater for GPT-family models, potentially biasing results and limiting apples-to-apples comparisons. Reproducibility is further hindered by confidentiality of internal artifacts and logs: we cannot release the full corpora. Future work includes leveraging production interaction logs (inputs, outputs, tool calls, actions) to mine candidate rules and hard examples that continuously enrich the policy tree and prompts, and implementing adaptive policy regeneration that automatically re-runs POLICY-TREE-GEN and POLICY-AS-PROMPT-GEN whenever PRDs, TDDs, code, or schemas undergo significant changes, with versioning and regression gating to mitigate drift. Also, prompt optimisations like Kumar et al. (2025) can be applied to POLICY-AS-PROMPT.
Appendix A Technical Appendices and Supplementary Material
A.1 POLICY-TREE-GEN PROMPTS
LLM Prompt: Pass 1 — Parse Document & Classify Security Instructions Role & Goal You are a document analysis expert specializing in data security for LLM-powered applications. Read a document describing an LLM-based system, infer its logical structure, and extract only those instructions relevant to data and guardrail security. Important: Ignore all specific examples in this step. Focus strictly on the instructions themselves. Input <Document text> Task Return a single JSON object representing the document’s hierarchy. For each extracted instruction, you must include: The exact, verbatim source_span copied from the document. A category chosen from 4 CATEGORIES. Valid Categories & Guidance VALINP: Defines what the system should accept (in-domain, topical, acceptable inputs). Exclude: outputs, rejection rules, metadata, formatting, or negative conditions. VALOUT: Defines correct system responses (including handling invalid/unsupported/out-of-domain inputs). Exclude: malformed prompts, format/schema rules, valid inputs, or outputs. INVALINP: Defines what the system must reject (invalid, unsupported, or out-of-domain inputs). Include: logic/guardrails for blocking inputs. Exclude: formatting, valid inputs, or outputs. INVALOUT: Defines incorrect or forbidden system responses (hallucinations, policy violations, leaking PII, misleading answers). Include: logic/guardrails for blocking outputs. Exclude: formats, constraints, valid outputs, or inputs. Exclusion Rule: If an instruction does not clearly fit one of the valid categories above, omit it from the output. Required JSON Shape Root key: "document_structure" (list). Each element: "topic": Heading "content": Instructions with source_span, category, sub_instructions "children": Sub-topics (same format)
Figure 2: Prompt for hierarchical parsing in POLICY-TREE-GEN
LLM Prompt: Pass 2 — Extract & Classify Policy Examples Role & Goal You are a policy expert. For a given policy instruction, find all relevant examples that already exist in the provided document. Do not invent or generate new examples. Inputs POLICY INSTRUCTION:
"{instruction_text}"
{document_text}
Task Return a single JSON object with a top-level key "examples" containing a list of example objects found in the document that are relevant to the policy instruction. Constraints No generation: Only extract examples that already exist in the document. Include every relevant example; omit non-relevant content. Example Object Schema Each example object must include: "example_text" — The text of the example. "source_span" — The exact, verbatim sentence copied from the document. "polarity" — Positive or Negative: positive: the example follows the instruction. negative: the example is a counter-example. Empty Result Handling If no relevant examples are found, return:
{"examples": []}
In-Domain Example (context only)
Figure 3: Prompt for extracting existing policy examples in POLICY-TREE-GEN
A.2 POLICY-AS-PROMPT OUTPUT
LLM Prompt: Policy Compliance Classification Task You are a meticulous security and policy compliance analyst. Your sole responsibility is to analyze a user’s input and determine if it complies with a strict set of predefined rules. Policy Document and Classification Reference Data:
Rules for VALID Inputs It is acceptable to automate the analysis of HR support tickets only if the content is free of personal data. Example of a VALID anonymized HR support ticket: "Request: I am unable to access the HR portal to update my benefits information. Error: System shows ’Access Denied’ despite using correct login. Resolution attempted: Cleared cache, tried on a different browser. Request: Please escalate to IT for resolution." Rules for INVALID Inputs Including any personal data in AI analysis or outputs without prior automated deletion or anonymization is a direct violation of the policy. Examples of personal data that must be removed include: names, email addresses, employee IDs, government IDs, birth dates, home addresses, salary details, beneficiary information, etc. Example of an INVALID non-anonymized HR support ticket: "Request: John Smith (Employee ID: 45832) cannot access his HR portal. Email: john.smith@company.com Issue: Unable to update dependent details for his son born on 12/03/2020. Request: Please resolve urgently." Example of an INVALID non-HR ticket (out of scope): "Request: The office coffee machine on the 3rd floor is broken again. Issue: Display shows ’Water Pump Error.’ Request: Please arrange maintenance." Output Format: Return a single JSON object with the following keys: "classification": Either "VALID" or "INVALID". "reasoning": A one-sentence explanation referencing the rule followed or violated. Task: Analyze the following DOCUMENT and provide your classification: Your JSON response:
Figure 4: Policy-As-Prompt for HR Application Input Classifier
LLM Prompt: Policy Compliance Classification (Content-Focused) You are a security/policy analyst. Analyze the user’s query and decide if the content of the INPUT is allowed under the policy. Ignore tooling/functions; focus only on what data the INPUT contains and what the OUTPUT must contain. Policy — Allowed vs. Disallowed Content ALLOWED INPUT CONTENT (no personal data, no secrets) Alert metadata: alert/case IDs, timestamps, severity, source system, rule/signature names, tags. Technical indicators (IoCs): file hashes, public IPs, domains, URLs (no credentials), file/process names, registry keys, event codes. System/asset context: asset IDs, generic hostnames (not derived from real names), service/application names, environment (prod/stage/dev), non-personal ticket/channel IDs. Log details (sanitized): network flow summaries, HTTP metadata (without tokens/cookies), counts, time windows, error codes; fields may be pseudonymized or hashed. Org context (non-personal): role or team names (e.g., “SRE on-call”), runbook references, anonymized user placeholders (user_A, hashed IDs). Threat intel: malware family names, campaign/actor names, rule IDs, confidence scores. Attachments/screenshots only if redacted/pseudonymized so no personal data or secrets are visible. DISALLOWED INPUT CONTENT Personal data (PII): names, emails, phone numbers, home addresses, government IDs, birth dates, employee/customer IDs tied to individuals. Sensitive records: health, financial, or HR records; salary/beneficiary details; customer lists; resumes. Secrets/credentials: passwords, API keys, private keys, tokens, session IDs, cookies, JWTs, MFA codes, OAuth headers. Raw dumps containing PII/secrets: browser/localStorage, memory dumps, unredacted chat/email transcripts, authorization headers. Identifiers revealing a person: hostnames/usernames or filenames that embed real names (must be pseudonymized before submission). Out-of-scope facility issues (e.g., coffee machine maintenance). Output Contract Return a single JSON object: "classification": "VALID" if the INPUT only contains Allowed content; otherwise "INVALID". "reasoning": One concise sentence citing the specific rule followed/violated. Task: Analyze the following DOCUMENT and classify it: Your JSON response:
Figure 5: Content-First Policy-As-Prompt for Security Input Classification
A.3 Evaluation for POLICY-TREE-GEN
Table 4: Evaluation Metrics for POLICY-TREE-GEN
Application Model Det P (%) Micro-F1 (%) Span Exact (%) Token-F1 Substr (%) Emb Cos
HR O1 68.6 30.0 91.7 0.999 100.0 0.996
HR GPT-OSS 120B 42.1 9.4 0.0 0.910 87.5 0.905
HR Llama 405B 40.0 10.9 75.0 1.000 100.0 0.973
HR Claude 3.5 40.0 4.0 50.0 0.992 100.0 0.986
SecOps O1 26.9 22.6 85.7 0.987 100 0.958
SecOps GPT-OSS 120B 66.7 10.3 0.0 0.974 0.0 0.881
SecOps Llama 405B 40.0 9.8 0.0 0.842 100.0 0.790
SecOps Claude 3.5 100.0 5.4 0.0 0.818 100.0 0.782