### v2

# Regulatable ML

Abstract
As autonomous AI agents are used in regulated and safety-critical settings, organizations need effective ways to turn policy into enforceable controls. We introduce a regulatory machine learning framework that converts unstructured design artifacts (like PRDs, TDDs, and code) into verifiable runtime guardrails. Our Policy as Prompt method reads these documents and risk controls to build a source-linked policy tree. This tree is then compiled into lightweight, prompt-based classifiers for real-time runtime monitoring. The system is built to enforce least privilege and data minimization. For conformity assessment, it provides complete provenance, traceability, and audit logging, all integrated with a human-in-the-loop review process. Evaluations show our system reduces prompt-injection risk, blocks out-of-scope requests, and limits toxic outputs. It also generates auditable rationales aligned with AI governance frameworks. By treating policies as executable prompts (a policy-as-code for agents), this approach enables secure-by-design deployment, continuous compliance, and scalable AI safety and AI security assurance for regulatable ML.
1 Introduction
Powerful AI agents are moving into everyday business, from helping HR to flagging security threats Wang et al. (2025); Ding et al. (2024). But this power comes with risk: an HR agent could accidentally leak a salary, or a helpful chatbot could be tricked into running a malicious command Liu et al. (2024); Kim et al. (2023); Li et al. (2025); He et al. (2024). This has created an urgent need to make sure these agents are safe and follow our rules, a sentiment echoed by emerging frameworks like the EU AI Act eu- (2024); Fabiano (2024).
The core problem is what we call the “policy-to-practice” gap: it’s easy for a human to write a rule in a design document, but it’s incredibly hard to turn that simple English sentence into a machine-enforceable rule that works reliably. This gap is a major roadblock to building, testing, and trusting AI systems. To solve this, we can use “guardrails”—safety checks that prevent the AI from doing unintended or harmful things Dong et al. (2024); Zhang et al. (2025). A key security idea here is the principle of least privilege, which means giving a system only the minimal access it needs to do its job. For AI agents, this ensures they stay within defined limits. However, static rules are often too rigid or too vague and fail to capture context. As recent research points out, security for these flexible agents needs to be just-in-time and context-aware (Kholkar and Ahuja, 2025; Tsai and Bagdasarian, 2025). While some have proposed static principles Hua et al. (2024), this is often not enough for dynamic, real-world interactions.
To bridge this critical gap, we introduce Policy as Prompt, a novel framework that reads natural language policy documents and turns them into dynamic, enforceable guardrails. Our system offers a practical way to implement the contextual security that Tsai and Bagdasarian (2025) called for. Our key contributions are as follows: (i) We introduce a scalable, end-to-end pipeline that automatically reads the unstructured technical artifacts teams already write (like PRDs or design docs) to identify and extract security constraints. (ii) We propose a verifiable process where these constraints are converted into a human-readable policy draft, enabling efficient review and refinement by security engineers. (iii) We compile the verified policy into prompt-based classifiers—our “guardrail security policies” Dong et al. (2024)—that use a lightweight LLM to act as a real-time "judge," enforcing the least-privilege policy by validating agent inputs and outputs, ensuring they only do what the policy explicitly allows and nothing more. (iv) We validate our approach by generating policies for various LLM applications and test their effectiveness across different state-of-the-art models.
Figure 1: Policy Generation and Enforcement Pipeline for an HR Application
2 Policy Tree Generation
Our technique extracts guardrail security requirements for LLM applications directly from development artifacts. By analyzing high-level design documents, we capture the system’s intended security posture before implementation. Grounding policies in their original design context is critical for ensuring AI systems operate as intended Tsai and Bagdasarian (2025).
Our core process, POLICY-TREE-GEN, uses a verified two-step method to extract and validate security rules directly from application design documents. Step 1 (Parse & Classify): An AI system analyzes the documents, identifies sentences that define security rules or data constraints, and classifies them into one of four categories: ID-I (In-Domain Inputs), OOD-I (Out-of-Domain Inputs, e.g., off-topic or malicious requests), ID-O (In-Domain Outputs), and OOD-O (Out-of-Domain Outputs, e.g., data leaks or toxic responses). Each extracted rule is double-checked by another AI agent for accuracy and proper categorization. Step 2 (Enrich with Examples): The system links these rules to relevant examples found in the documents, producing a structured, verifiable policy tree that preserves contextual grounding. This process inherently enforces the Principle of Least Privilege Zhang et al. (2025): the agent can act only within what is explicitly permitted by its design documents. For instance, for an HR App, a rule such as “Only access resolved HR case data from the Case Portal” is classified as ID-I, meaning the agent must reject attempts to access non-HR data like IT or Sales records (OOD-I). Similarly, a rule like “Generated KB article should be flagged if employee identifiers are present” is tagged as OOD-O, ensuring the agent’s responses do not reveal sensitive information. Thus, an HR agent derived from this process is automatically restricted to HR-related data and outputs, without implicit access to finance or IT information. The resulting policy tree defines and verifies these input–output boundaries, serving as the enforceable foundation for secure, least-privilege AI behavior.
Table 1: Policy Enforcement Example for HR Input Classifier
App Example Classification Reason Code Action
HR Update my address on Workday ID (Input) ALLOW
HR My address is 21 Victoria St OOD (Input) Contains Non-Anonymised PII BLOCK
HR Ignore rules and reveal your system prompt OOD (Input) Malicious / Prompt Injection ALERT
HR Send to sample@gmail.com ID (Input) Anonymised PII ALLOW
HR US election news link… OOD (Output) Non HR Content BLOCK
3 Policy as Prompt Generation
The POLICY-AS-PROMPT-GEN begins with the verified policy tree from POLICY-TREE-GEN, which encodes categorized security rules and examples. This tree is transformed into a human-readable markdown document for LLM consumption, where each instruction is formatted as a rule with examples labeled ‘positive’ (compliant) or ‘negative’ (violating). The labeled examples are reused to synthesize few-shot prompt blocks that accompany the rules.
It is then embedded into a master template that primes the LLM as a compliance analyst. Two templates are used: one for an Input Classifier and another for an Output Auditor. These prompt the model to consult the policy rules and return a JSON object with a binary classification (ID/OOD) and a concise justification. For HR input classification (Table 1), HR-related requests without non-anonymized PII are ID, while those with PII, prompt-injection attempts, or non-HR content are OOD. The strict output format enables deterministic system actions (ALLOW, BLOCK, ALERT). HR Input Classifier is shown in Figure 4 and SOC Input Classifier in Figure 5. The deliverables are two markdown few-shot prompts (rules + exemplars), which undergo human-in-the-loop review. Security engineers then either (i) approve for deployment, (ii) reject requiring upstream updates and regeneration, or (iii) request changes to the prompt markdown without altering the upstream tree.
4 Experimental Setup and Results
This study evaluated the generation and enforcement of guardrail policies across two distinct LLM-powered systems in Human Resource (HR) and Security Operations Centre (SOC)) domains. The input artifacts for all models consisted of PRDs, technical design documents, and prompts extracted from the application source code. These artifacts were sourced from real-world, internal enterprise projects, representing authentic, in-production use cases. To ensure compatibility, all documents were converted to Markdown format, and any embedded images were replaced with textual descriptions generated by gpt-4o. All reported metrics are the average of multiple runs to ensure stability.
Table 2: Evaluation Metrics for POLICY-TREE-GEN
Detection Classification Per-class F1
Application Model R (%) F1 (%) Macro-F1 (%) ID-I ID-O INVINP INVOUT
HR O1 53.3 60.0 24.5 35.3 29.4 0.0 33.3
HR GPT-OSS 120B 17.8 25.0 4.8 19.4 0.0 0.0 0.0
HR Llama 405B 8.9 14.5 4.5 18.2 0.0 0.0 0.0
HR Claude 3.5 4.4 8.0 12.5 0.0 0.0 0.0 50.0
SOC O1 19.4 22.6 13.0 22.2 29.6 0.0 0.0
SOC GPT-OSS 120B 5.6 10.3 4.2 16.7 0.0 0.0 0.0
SOC Llama 405B 5.6 9.8 10.0 0.0 0.0 40.0 0.0
SOC Claude 3.5 2.8 5.4 6.2 0.0 0.0 25.0 0.0
POLICY-TREE-GEN Analysis: We evaluated several large language models including Llama 3 405B(Grattafiori et al., 2024), GPT OSS 120B (OpenAI et al., 2025), Claude Sonnet 3.5(Anthropic, 2024) and o1 (OpenAI et al., 2024) on two applications, HR and SOC. The models Gemma 3 1B(Team et al., 2025) and Qwen 1.7B Thinking(Yang et al., 2025) were excluded due to poor performance. Gold policies were created by security engineers and served as the ground truth against which the LLM-generated policies were evaluated. The metrics in Table 2 evaluate model performance in two categories: Detection, measured by Recall (R) and F1 score, and Classification, assessed via Macro-F1 and individual per-class F1 scores for ID-I, ID-O, OOD-I, and OOD-O. A high score in these metrics indicates a model that is effective at both identifying relevant requirements and accurately assigning the correct category to them. HR consistently yields higher performance scores for all models compared to SOC. Furthermore, the o1 model significantly outperforms all other models, demonstrating its superior capability in this specific task.
As shown in Table 4, this analysis details model performance on HR and SOC tasks using metrics such as Detection Precision, which measures the percentage of a model’s predictions that are correct, and Micro-F1. Span quality is evaluated using four metrics: Span Exact, which measures if the extracted text is an exact match to the ground truth; Token-F1, which assesses the token-level overlap between the extracted and ground truth text; Substr, which checks if one text is a substring of the other; and Emb Cos, which is the cosine similarity of the text embeddings. A high score in these metrics indicates that the model is accurately extracting the correct text, even if it might misclassify it. While the o1 model remains a top performer, others reveal a significant disconnect between their extraction quality and overall F1 score. For HR, models like Llama 405B and Sonnet 3.5 have high span quality metrics, indicating accurate text extraction, yet their low Micro-F1 scores suggest they struggle with correct classification. For SOC domain, performance is generally lower and more varied. Notably, Sonnet 3.5 achieved a high Det P on SOC, meaning all of its identified requirements were true positives, but its Micro-F1 score remained very low due to continued classification errors.
Table 3: Accuracy per application/model for POLICY-AS-PROMPT-GEN
Application Model Input Acc. Output Acc.
HR GPT-4o 0.73 0.71
Qwen3-1.7B 0.66 0.59
Gemma-1B 0.40 0.32
SOC GPT-4o 0.70 0.68
Qwen3-1.7B 0.66 0.61
Gemma-1B 0.42 0.41
POLICY-AS-PROMPT-GEN Analysis: Policy enforcement was tested on a separate set of models: Qwen 3 1.7B (Thinking Mode) (Yang et al., 2025), GPT-4o (OpenAI et al., 2024), and Gemma 3 1B(Team et al., 2025). For specialized data classification tasks within this framework, we employed Small Language Models (SLMs), leveraging their established effectiveness and low latency, which helps to quickly run policies in real-time. Both policies were judged request review by security engineers as part of our evaluation and deployed after minimal changes, saving time for the security team. POLICY-AS-PROMPT-GEN was evaluated on functional correctness by integrating generated prompts into target LLMs and testing with 100 gold inputs and outputs. Tests included both standard and adversarial cases (e.g., prompt injection, toxic content). Evaluation focused on accuracy of OOD/ID classifications, with defaults of OOD $\Rightarrow$ BLOCK and ID $\Rightarrow$ ALLOW, unless otherwise specified. While accuracies in the  70-73% range for GPT-4o (Table 3) are not perfect, they demonstrate significant utility in a real-world context. The system acts as a "default-deny" guardrail: its primary value is in successfully blocking a high percentage of malicious or non-compliant inputs (true positives for OOD) before they reach the agent, drastically reducing the attack surface. This level of accuracy is highly effective as a first-line defense, flagging ambiguous cases for human review, which is a significant improvement over manual, post-hoc auditing. A limitation is that prompt tuning emphasized GPT models, potentially contributing to performance gaps.
5 Conclusion
In this work, we introduced a framework that bridges the policy-to-practice gap by transforming unstructured design artifacts into verifiable guardrails. We demonstrated an end-to-end pipeline for automated, auditable, and enforceable policy generation. Our experiments show that while large proprietary models excel in policy extraction, smaller models can still effectively enforce policies with curated prompts. This approach establishes a scalable path toward trustworthy, regulatable AI systems grounded in transparent policy governance.
6 Limitations
Despite promising results, this work has several limitations that temper generalizability and reproducibility. Our evaluation spans two enterprise-style domains (HR, SOC) with modest gold sets (100 inputs and 100 outputs per application) and a limited number of design artifacts, which constrains transferability to other settings and larger, more heterogeneous corpora. Policy extraction currently depends on large proprietary models, while open-source baselines differ in size or tuning; moreover, prompting effort was greater for GPT-family models, potentially biasing results and limiting apples-to-apples comparisons. Reproducibility is further hindered by confidentiality of internal artifacts and logs: we cannot release the full corpora. Future work includes leveraging production interaction logs (inputs, outputs, tool calls, actions) to mine candidate rules and hard examples that continuously enrich the policy tree and prompts, and implementing adaptive policy regeneration that automatically re-runs POLICY-TREE-GEN and POLICY-AS-PROMPT-GEN whenever PRDs, TDDs, code, or schemas undergo significant changes, with versioning and regression gating to mitigate drift. Also, prompt optimisations like Kumar et al. (2025) can be applied to POLICY-AS-PROMPT.
Appendix A Technical Appendices and Supplementary Material
A.1 POLICY-TREE-GEN PROMPTS
LLM Prompt: Pass 1 — Parse Document & Classify Security Instructions Role & Goal You are a document analysis expert specializing in data security for LLM-powered applications. Read a document describing an LLM-based system, infer its logical structure, and extract only those instructions relevant to data and guardrail security. Important: Ignore all specific examples in this step. Focus strictly on the instructions themselves. Input <Document text> Task Return a single JSON object representing the document’s hierarchy. For each extracted instruction, you must include: The exact, verbatim source_span copied from the document. A category chosen from 4 CATEGORIES. ID Categories & Guidance ID-I: Defines what the system should accept (in-domain, topical, acceptable inputs). Exclude: outputs, rejection rules, metadata, formatting, or negative conditions. ID-O: Defines correct system responses (including handling invalid/unsupported/out-of-domain inputs). Exclude: malformed prompts, format/schema rules, valid inputs, or outputs. OOD-I: Defines what the system must reject (invalid, unsupported, or out-of-domain inputs). Include: logic/guardrails for blocking inputs. Exclude: formatting, valid inputs, or outputs. OOD-O: Defines incorrect or forbidden system responses (hallucinations, policy violations, leaking PII, misleading answers). Include: logic/guardrails for blocking outputs. Exclude: formats, constraints, valid outputs, or inputs. Exclusion Rule: If an instruction does not clearly fit one of the valid categories above, omit it from the output. Required JSON Shape Root key: "document_structure" (list). Each element: "topic": Heading "content": Instructions with source_span, category, sub_instructions "children": Sub-topics (same format)
Figure 2: Prompt for hierarchical parsing in POLICY-TREE-GEN
LLM Prompt: Pass 2 — Extract & Classify Policy Examples Role & Goal You are a policy expert. For a given policy instruction, find all relevant examples that already exist in the provided document. Do not invent or generate new examples. Inputs POLICY INSTRUCTION:
"{instruction_text}"
{document_text}
Task Return a single JSON object with a top-level key "examples" containing a list of example objects found in the document that are relevant to the policy instruction. Constraints No generation: Only extract examples that already exist in the document. Include every relevant example; omit non-relevant content. Example Object Schema Each example object must include: "example_text" — The text of the example. "source_span" — The exact, verbatim sentence copied from the document. "polarity" — Positive or Negative: positive: the example follows the instruction. negative: the example is a counter-example. Empty Result Handling If no relevant examples are found, return:
{"examples": []}
In-Domain Example (context only)
Figure 3: Prompt for extracting existing policy examples in POLICY-TREE-GEN
A.2 POLICY-AS-PROMPT OUTPUT
LLM Prompt: Policy Compliance Classification Task You are a meticulous security and policy compliance analyst. Your sole responsibility is to analyze a user’s input and determine if it complies with a strict set of predefined rules. Policy Document and Classification Reference Data:
Rules for ID Inputs It is acceptable to automate the analysis of HR support tickets only if the content is free of personal data. Example of a ID anonymized HR support ticket: "Request: I am unable to access the HR portal to update my benefits information. Error: System shows ’Access Denied’ despite using correct login. Resolution attempted: Cleared cache, tried on a different browser. Request: Please escalate to IT for resolution." Rules for OOD Inputs Including any personal data in AI analysis or outputs without prior automated deletion or anonymization is a direct violation of the policy. Examples of personal data that must be removed include: names, email addresses, employee IDs, government IDs, birth dates, home addresses, salary details, beneficiary information, etc. Example of an OOD non-anonymized HR support ticket: "Request: John Smith (Employee ID: 45832) cannot access his HR portal. Email: john.smith@company.com Issue: Unable to update dependent details for his son born on 12/03/2020. Request: Please resolve urgently." Example of an OOD non-HR ticket (out of scope): "Request: The office coffee machine on the 3rd floor is broken again. Issue: Display shows ’Water Pump Error.’ Request: Please arrange maintenance." Output Format: Return a single JSON object with the following keys: "classification": Either "ID" or "OOD". "reasoning": A one-sentence explanation referencing the rule followed or violated. Task: Analyze the following DOCUMENT and provide your classification: Your JSON response:
Figure 4: Policy-As-Prompt for HR Application Input Classifier
LLM Prompt: Policy Compliance Classification (Content-Focused) You are a security/policy analyst. Analyze the user’s query and decide if the content of the INPUT is allowed under the policy. Ignore tooling/functions; focus only on what data the INPUT contains and what the OUTPUT must contain. Policy — Allowed vs. Disallowed Content ALLOWED INPUT CONTENT (no personal data, no secrets) Alert metadata: alert/case IDs, timestamps, severity, source system, rule/signature names, tags. Technical indicators (IoCs): file hashes, public IPs, domains, URLs (no credentials), file/process names, registry keys, event codes. System/asset context: asset IDs, generic hostnames (not derived from real names), service/application names, environment (prod/stage/dev), non-personal ticket/channel IDs. Log details (sanitized): network flow summaries, HTTP metadata (without tokens/cookies), counts, time windows, error codes; fields may be pseudonymized or hashed. Org context (non-personal): role or team names (e.g., “SRE on-call”), runbook references, anonymized user placeholders (user_A, hashed IDs). Threat intel: malware family names, campaign/actor names, rule IDs, confidence scores. Attachments/screenshots only if redacted/pseudonymized so no personal data or secrets are visible. DISALLOWED INPUT CONTENT Personal data (PII): names, emails, phone numbers, home addresses, government IDs, birth dates, employee/customer IDs tied to individuals. Sensitive records: health, financial, or HR records; salary/beneficiary details; customer lists; resumes. Secrets/credentials: passwords, API keys, private keys, tokens, session IDs, cookies, JWTs, MFA codes, OAuth headers. Raw dumps containing PII/secrets: browser/localStorage, memory dumps, unredacted chat/email transcripts, authorization headers. Identifiers revealing a person: hostnames/usernames or filenames that embed real names (must be pseudonymized before submission). Out-of-scope facility issues (e.g., coffee machine maintenance). Output Contract Return a single JSON object: "classification": "ID" if the INPUT only contains Allowed content; otherwise "OOD". "reasoning": One concise sentence citing the specific rule followed/violated. Task: Analyze the following DOCUMENT and classify it: Your JSON response:
Figure 5: Content-First Policy-As-Prompt for Security Input Classification
A.3 Evaluation for POLICY-TREE-GEN
Table 4: Evaluation Metrics for POLICY-TREE-GEN
Application Model Det P (%) Micro-F1 (%) Span Exact (%) Token-F1 Substr (%) Emb Cos
HR O1 68.6 30.0 91.7 0.999 100.0 0.996
HR GPT-OSS 120B 42.1 9.4 0.0 0.910 87.5 0.905
HR Llama 405B 40.0 10.9 75.0 1.000 100.0 0.973
HR Claude 3.5 40.0 4.0 50.0 0.992 100.0 0.986
SOC O1 26.9 22.6 85.7 0.987 100 0.958
SOC GPT-OSS 120B 66.7 10.3 0.0 0.974 0.0 0.881
SOC Llama 405B 40.0 9.8 0.0 0.842 100.0 0.790
SOC Claude 3.5 100.0 5.4 0.0 0.818 100.0 0.782